OpenStudy (anonymous):

When the quadratic equation y = (x − 3)(4x + 2) is written in standard form, what is the value of the coefficient "b?" 1 4 −10 −6

OpenStudy (anonymous):

y=4x^2-12x+2x-6 y=4x^2-10x-6 y=ax^2+bx+c b=-10